Zinc reacts with hydrochloric acid according to th
Answer #1

Zn mass = ( 21/100) x 8.45 g = 1.7745 g

Zn moles = mass / atomic mass of Zn

            = ( 1.7745 g) / ( 65.38 g/mol)

                    = 0.02714

HCl moles needed = 2 x Zn moles   ( as per balanced eq)

                   = 2 x 0.02714 = 0.0543

HCl molarity given 3.5 M

Molarity = moles / volume

3.5 = 0.0543 / ( volume)

volume = 0.0155 L = 15.5 ml

Thus volume of HCl needed = 15.5 ml